### v1

# Representation Learning for Distributional Perturbation Extrapolation

Abstract
We consider the problem of modelling the effects of unseen perturbations such as gene knockdowns or drug combinations on low-level measurements such as RNA sequencing data. Specifically, given data collected under some perturbations, we aim to predict the distribution of measurements for new perturbations. To address this challenging extrapolation task, we posit that perturbations act additively in a suitable, unknown embedding space. More precisely, we formulate the generative process underlying the observed data as a latent variable model, in which perturbations amount to mean shifts in latent space and can be combined additively. Unlike previous work, we prove that, given sufficiently diverse training perturbations, the representation and perturbation effects are identifiable up to affine transformation, and use this to characterize the class of unseen perturbations for which we obtain extrapolation guarantees. To estimate the model from data, we propose a new method, the perturbation distribution autoencoder (PDAE), which is trained by maximising the distributional similarity between true and predicted perturbation distributions. The trained model can then be used to predict previously unseen perturbation distributions. Empirical evidence suggests that PDAE compares favourably to existing methods and baselines at predicting the effects of unseen perturbations.
1 Introduction
Due to technological progress, large-scale perturbation data is becoming more abundant across several scientific fields. This is particularly the case for single-cell biology, where advancements in gene editing, sequencing, and mass spectrometry have led to the collection of vast transcriptomic or proteomic databases for various drug and gene perturbations (Wang et al., 2009; Jinek et al., 2012; Weinstein et al., 2013; Dixit et al., 2016; Norman et al., 2019). However, the exponential number of possible combinations of perturbations renders exhaustive experimentation prohibitive. Observations are thus typically only available for a subset of perturbations of interest, e.g., some single and double gene knockdowns or certain dosages of drugs. Having models capable of extrapolating to unseen combinations of perturbations, e.g., new multi-gene knockdowns or dosage combinations, would be highly desirable.
Prior Work.
Several recent works advocate for the use of machine learning for biological perturbation modelling (Lopez et al., 2018, 2023), often with the goal of generalizing to new cell types (Lotfollahi et al., 2019; Bunne et al., 2023), unseen combinations of perturbations (Lotfollahi et al., 2023; Bereket & Karaletsos, 2023), or entirely new perturbations by leveraging the molecular structure of the involved compounds (Yu & Welch, 2022; Hetzel et al., 2022; Qi et al., 2024) or prior knowledge about gene-gene interactions (Roohani et al., 2024; Kamimoto et al., 2023). A common theme is the use of nonlinear representation learning techniques such as autoencoders (Rumelhart et al., 1986; Hinton & Salakhutdinov, 2006; Kingma & Welling, 2014) to embed observations in a latent space, in which the effects of perturbations are assumed to take on a simpler (e.g., additive) form. However, existing studies are purely empirical and lack theoretical underpinning. Despite promising results, the capabilities and fundamental limitations of existing methods thus remain poorly understood.
Overview and Contributions.
In this work, we present a principled, theoretically-grounded approach for perturbation extrapolation. Given the unpaired nature of the available non-i.i.d. data (there are no correspondences between observations from different domains since each cell is only measured under one experimental condition) we consider the task of predicting population-level effects of perturbations, which we formalize as a distributional regression problem (§ 2). We then postulate a generative model (§ 3) which, similar to prior works, assumes that perturbations act as mean shifts in a suitable latent space, see Fig. 1 for an overview. For the case of a deterministic decoder, we analyse this model class theoretically (§ 4), proving that, under mild assumptions, the latent representation and the relative training perturbation effects are identifiable up to affine transformation if the available environments are sufficiently diverse (Thm. 4.1). This result implies extrapolation guarantees for unseen perturbations that can be expressed as linear combinations of training perturbations (Thm. 4.6). Based on these insights, we devise an autoencoder-based estimation method (§ 5) which draws inspiration from distributional principal autoencoders (Shen & Meinshausen, 2024a) and uses the energy score (Gneiting & Raftery, 2007) to assess the distributional similarity between predicted and ground-truth perturbation data. In simulations (§ 7), our approach compares favourably to the compositional perturbation autoencoder (CPA; Lotfollahi et al., 2023) and several baselines in terms of mean prediction and distributional fit.
Notation.
[…]
(a) Simulated Multi-Domain Perturbation Data
[…]
(b) Graphical Model Representation
Figure 1: Task Description and Example Following the Assumed Data Generating Process. (a) During training, we are given $M\!=\!5$ training data sets in observation space (right, contour plots in grey), each of which is generated under a known combination of $K\!=\!3$ elementary perturbations. The corresponding (training) perturbation labels $\mathbf{l}_{e}$ are shown below the plots. During testing, we are given a new perturbation label and the task is to predict the corresponding distribution in observation space (right, blue and orange). We tackle this task by assuming that the effect of perturbations is linear additive in a suitable latent space (left). Both plots show kernel density estimates of the distributions. (b) For each experiment or environment $e\in[M]_{0}$, the corresponding dataset comprises a perturbation label $\bm{l}_{e}$ and $N_{e}$ observations $\bm{x}_{e,i}$. Perturbations are assumed to act as mean-shifts on a latent basal state, $\mathbf{z}^{\text{pert}}_{e,i}=\mathbf{z}^{\text{base}}_{e,i}+\mathbf{W}\mathbf{l}_{e}$ for an unknown perturbation matrix $\mathbf{W}$. An (unknown) stochastic nonlinear decoder with noise $\bm{\varepsilon}_{e,i}$ then yields the observed $\mathbf{x}_{e,i}=\mathbf{f}(\mathbf{z}^{\mathrm{pert}}_{e,i},\bm{\varepsilon}_{e,i})$. In the example in (a), $\mathbf{f}$ is constant in its second argument. Shaded and white nodes indicate observed and unobserved/latent variables, respectively.
2 Problem Setting: Distributional Perturbation Extrapolation
Observables.
Let $\mathbf{x}\in\mathbb{R}^{{d_{X}}}$ be an observation (e.g., omics data) that is obtained under one of several possible experimental conditions, and let $\mathbf{X}$ be the corresponding random vector. We model these conditions as combinations of $K$ elementary perturbations, each of which we assume can be encoded by a real number. Further, let $\mathbf{l}\in\mathbb{R}^{K}$ be a vector of perturbation labels that indicates if, or how much of, each perturbation was applied before collecting $\mathbf{x}$.
Example 2.1 (Gene perturbations).
For data arising from gene knockouts, a perturbation can be represented by a binary $\mathbf{l}\in\{0,1\}^{K}$, where $K$ denotes the number of potential targets and $l_{k}=1$ if and only if target $k$ was subject to a knockout experiment. For example, $\mathbf{l}=(1,0,1,0,0)$ indicates a multi-gene knockout on targets one and three.
Example 2.2 (Drug perturbations).
For data arising from applying varying amounts of $K$ different drugs, perturbations can be represented by continuous, non-negative labels $\mathbf{l}\in\mathbb{R}_{+}^{K}$, where $l_{k}$ indicates the (relative or absolute) amount of drug $k$ that was administered.
Data.
We have access to $M+1$ experimental datasets $\mathcal{D}_{0},\mathcal{D}_{1},...,\mathcal{D}_{M}$, each comprising a sample of $N_{e}$ observations $\mathbf{x}$ and a perturbation label $\mathbf{l}$, i.e., for all experiments or environments $e\in[M]_{0}$,
$\mathcal{D}_{e}=\left(\left(\mathbf{x}_{e,i}\right)_{i=1}^{N_{e}},\mathbf{l}_{e}\right).$ (2.1)
For all $e\in[M]_{0}$, we assume that $\mathbf{x}_{e,1},...,\mathbf{x}_{e,N_{e}}$ are i.i.d. realization from some underlying distribution $\mathbb{P}_{\mathbf{X}|\mathbf{l}_{e}}$.
Task.
Given data in the form of (2.1), we aim to predict the effects of new perturbations $\mathbf{l}_{\mathrm{test}}\not\in\{\mathbf{l}_{0},\mathbf{l}_{1},...,\mathbf{l}_{M}\}$ without observing any data from this condition (“zero-shot”). In particular, we are interested in the distribution over observations resulting from $\mathbf{l}_{\mathrm{test}}$. That is, we aim to leverage the training domains (2.1) to learn a map
$\mathbf{l}\mapsto\mathbb{P}_{\mathbf{X}|\mathbf{l}},$ (2.2)
which extrapolates beyond the training support of $\mathbf{l}$, i.e., the predictions should remain reliable for new inputs $\mathbf{l}_{\mathrm{test}}$.
Distributional vs. Mean Prediction.
Since (2.2) targets the full conditional distribution—rather than, say, the conditional mean $\mathbb{E}[\mathbf{X}|\mathbf{l}]$—it constitutes a (multi-variate) distributional regression task (Koenker & Bassett Jr, 1978; Koenker, 2005), also referred to as probabilistic forecasting (Gneiting & Raftery, 2007) or conditional generative modelling (Mirza, 2014; Sohn et al., 2015; Winkler et al., 2019). Therefore, we refer to the problem setting addressed in this paper as distributional perturbation extrapolation.
On Extrapolation.
Formally, extrapolation means that the value of the function (2.2) at $\mathbf{l}_{\mathrm{test}}$ is determined by its values on the training support $\{\mathbf{l}_{0},\mathbf{l}_{1},...,\mathbf{l}_{M}\}$. Intuitively, for this to be feasible without imposing very restrictive assumptions on the form of the mapping in (2.2), $\mathbf{l}_{\mathrm{test}}$ should be somehow related to the training perturbations $\mathbf{l}_{e}$. For example, given data resulting from individual perturbations, predict the effects of combinations thereof. This type of extrapolation to new combinations of inputs is also called compositional generalization (Lake et al., 2017; Goyal & Bengio, 2022). It is known to be challenging (Schott et al., 2022; Montero et al., 2021, 2022) and requires assumptions that sufficiently constrain the model class  (Wiedemer et al., 2024a, b; Lachapelle et al., 2023; Brady et al., 2023, 2025; Dong & Ma, 2022; Lippl & Stachenfeld, 2024).
3 Model: Perturbations as Mean Shifts in Latent Space
We now specify a generative process for the observed data in (2.1). In so doing, we aim to strike a balance between imposing sufficient structure on (2.2) to facilitate extrapolation, while remaining flexible enough to model the complicated, nonlinear effects which perturbations may have on the distribution of observations. Similar to prior work (Lotfollahi et al., 2023; Bereket & Karaletsos, 2023), we model the effect of perturbations as mean shifts in a latent space with perturbation-relevant latent variables $\mathbf{z}\in\mathbb{R}^{{d_{Z}}}$, [^1] Xu et al. (2024) propose tests for this assumption, given suitable data from individual and combinations of perturbations. which are related to the observations $\mathbf{x}$ via a nonlinear (stochastic) mixing function or generator $\mathbf{f}$. The full generative process amounts to a hierarchical latent variable model, which additionally contains noise variables ${\bm{\varepsilon}}$ that capture other variation underlying the observations $\mathbf{x}$, and which is represented as a graphical model in Fig. 1(b). Specifically, we posit for all $e\in[M]_{0}\cup\{\mathrm{test}\}$ and all $i\in[N_{e}]$:
$\mathbf{Z}^{\mathrm{base}}_{e,i}$ $\sim\mathbb{P}_{\mathbf{Z}},$ (3.1)
$\mathbf{Z}^{\mathrm{pert}}_{e,i}$ $:=\mathbf{Z}^{\mathrm{base}}_{e,i}+\mathbf{W}\mathbf{l}_{e},$ (3.2)
${\bm{\varepsilon}}_{e,i}$ $\sim\mathbb{Q}_{\bm{\varepsilon}},$ (3.3)
$\mathbf{X}_{e,i}$ $:=\mathbf{f}\left(\mathbf{Z}^{\mathrm{pert}}_{e,i},{\bm{\varepsilon}}_{e,i}\right),$ (3.4)
where $(\mathbf{Z}^{\mathrm{base}}_{e,i})_{e\in[M]_{0},i\in[N_{e}]}$ are independent and identically distributed (i.i.d.) according to $\mathbb{P}_{\mathbf{Z}}$, and $({\bm{\varepsilon}}^{\mathrm{base}}_{e,i})_{e\in[M]_{0},i\in[N_{e}]}$ are i.i.d. according to $\mathbb{Q}_{\bm{\varepsilon}}$ and jointly independent of $(\mathbf{Z}^{\mathrm{base}}_{e,i})_{e\in[M]_{0},i\in[N_{e}]}$.
The basal state $\mathbf{Z}^{\mathrm{base}}$ in (3.1) describes the unperturbed state of latent variables, which can, in principle, be affected by perturbations, and is distributed according to a base distribution $\mathbb{P}_{\mathbf{Z}}$. The perturbation matrix $\mathbf{W}\in\mathbb{R}^{{d_{Z}}\times K}$ in (3.2) captures the effect of the $K$ elementary perturbations encoded in $\mathbf{l}$ on the latents and turns $\mathbf{Z}^{\mathrm{base}}$ into perturbed latents $\mathbf{Z}^{\mathrm{pert}}$. Since the same perturbation $\mathbf{l}_{e}$ is applied for all $i\in[N_{e}]$, all intra-dataset variability in $\mathbf{Z}^{\mathrm{pert}}$ is due to $\mathbb{P}_{\mathbf{Z}}$. The noise variables ${\bm{\varepsilon}}\in\mathbb{R}^{{d_{\epsilon}}}$ in (3.3) capture all other variation in the observed data that is unaffected by perturbations. W.l.o.g., we assume that it is distributed according to a fixed, uninformative distribution $\mathbb{Q}_{\bm{\varepsilon}}$ such as a standard isotropic Gaussian. The noise serves as an additional input to the (stochastic) mixing function or generator $\mathbf{f}:\mathbb{R}^{d_{Z}}\times\mathbb{R}^{d_{\epsilon}}\to\mathbb{R}^{d_{X}}$ in (3.4), which produces observations for the perturbed latent. This generative model allows us to capture any conditional distribution $\mathbb{P}_{\mathbf{X}|\mathbf{Z}=\mathbf{z}}$ (see, e.g., Peters et al., 2017, Prop. 7.1) and is more flexible than relying on parametric assumptions, such as a Gaussian (Lotfollahi et al., 2023) or negative Binomial (Bereket & Karaletsos, 2023; Lopez et al., 2018, 2023) likelihood parametrised by a deterministic decoder.
For a given $e$ and $\mathbf{l}_{e}$, the generative process in (3.1)–(3.4) induces a distribution over observations $\mathbf{x}$, which we denote by $\mathbb{P}_{e}$ or $\mathbb{P}_{\mathbf{X}|\mathbf{l}_{e}}$, defined as the push-forward of $\mathbb{P}_{\mathbf{Z}}$ and $\mathbb{Q}_{\bm{\varepsilon}}$ through (3.2) and (3.4), such that:
[…]
4 Theory: Identifiability and Extrapolation Guarantees
In this section, we present our theoretical analysis for the model class introduced in § 3. Here, we assume that the decoder is deterministic, so that we can write $\mathbf{f}:\mathbb{R}^{d_{Z}}\to\mathbb{R}^{d_{X}}$, see Remark 4.5 below for further discussion.
4.1 Identifiability
We first study identifiability, that is, the question under what assumptions and up to what ambiguities certain parts of the postulated generative process can be provably recovered assuming access to the full distributions. As established by the following results, our model class is identifiable up to affine transformation, provided that the training perturbations are sufficiently diverse, the dimension of the latent space is known and some additional technical assumptions hold.
Theorem 4.1 (Affine identifiability for Gaussian latents).
For $M\in\mathbb{Z}_{\geq 0}$, let $\mathbf{l}_{0},...,\mathbf{l}_{M}\in\mathbb{R}^{K}$ be ${M{+}1}$ perturbation labels. Let $\mathbf{f},\widetilde{\mathbf{f}}:\mathbb{R}^{{d_{Z}}}\to\mathbb{R}^{d_{X}}$, $\mathbf{W},\widetilde{\mathbf{W}}\in\mathbb{R}^{{d_{Z}}\times K}$, and $\mathbb{P},\widetilde{\mathbb{P}}$ be distributions on $\mathbb{R}^{{d_{Z}}}$ such that the models $(\mathbf{f},\mathbf{W},\mathbb{P})$ and $(\widetilde{\mathbf{f}},\widetilde{\mathbf{W}},\widetilde{\mathbb{P}})$ induce the same observed distributions, i.e.,
[…]
Assume further that:
  1. (i)
    [invertibility] $\mathbf{f}$ and $\widetilde{\mathbf{f}}$ are $C^{2}$-diffeomorphisms onto their respective images;
  2. (ii)
[…]
  3. (iii)
    [sufficient diversity] the matrix $\widetilde{\mathbf{W}}\mathbf{L}\in\mathbb{R}^{{d_{Z}}\times M}$, where $\mathbf{L}\in\mathbb{R}^{K\times M}$ is the matrix with columns $(\mathbf{l}_{1}-\mathbf{l}_{0})$, …, $(\mathbf{l}_{M}-\mathbf{l}_{0})$, has full row rank, i.e., $\mathrm{rank}(\widetilde{\mathbf{W}}\mathbf{L})={d_{Z}}$, or the same holds for $\mathbf{W}\mathbf{L}$.
Then the latent representation and the effects of the observed perturbation combinations relative to $\mathbf{l}_{0}$ (as captured by $\mathbf{W}\mathbf{L}$) are identifiable up to affine transformation in the following sense:
$\forall\mathbf{z}\in\mathbb{R}^{d_{Z}}:\qquad\widetilde{\mathbf{f}}^{-1}\circ\mathbf{f}(\mathbf{z})$ $=\mathbf{A}\mathbf{z}+\mathbf{b},$ (4.2)
$\widetilde{\mathbf{W}}\mathbf{L}$ $=\mathbf{A}\mathbf{W}\mathbf{L},$ (4.3)
where $\mathbf{A}:=\widetilde{\bm{\Sigma}}^{\frac{1}{2}}\bm{\Sigma}^{-\frac{1}{2}}$ and $\mathbf{b}:=\widetilde{\bm{\mu}}-\mathbf{A}\bm{\mu}+(\widetilde{\mathbf{W}}-\mathbf{A}\mathbf{W})\mathbf{l}_{0}$.
Corollary 4.2 (Affine recovery of the perturbation matrix).
If, in addition to the assumptions of Thm. 4.1, $\mathbf{L}\in\mathbb{R}^{K\times M}$ has full row rank (i.e., $\mathrm{rank}(\mathbf{L})=K\leq M$), then the perturbation matrix $\mathbf{W}$ is identifiable up to affine transformation in the sense that
$\widetilde{\mathbf{W}}=\mathbf{A}\mathbf{W},$ (4.4)
for $\mathbf{A}:=\widetilde{\bm{\Sigma}}^{\frac{1}{2}}\bm{\Sigma}^{-\frac{1}{2}}$. In this case, the expression for $\mathbf{b}$ in (4.2) simplifies to $\mathbf{b}=\widetilde{\bm{\mu}}-\mathbf{A}\bm{\mu}$.
The proofs of Thms. 4.1 and 4.2 are provided in §§ A.1 and A.2.
Discussion.
Thm. 4.1 can be interpreted as follows. Fix a set of perturbation labels $(\mathbf{l}_{0},...,\mathbf{l}_{M})$ and a data generating process parametrised by $(\mathbf{f},\mathbf{W},\bm{\mu},\bm{\Sigma})$. Then, for all $\widetilde{\bm{\mu}}$ and $\widetilde{\bm{\Sigma}}$, and for all $\widetilde{\mathbf{W}}$ such that (4.3) holds for $\mathbf{A}=\widetilde{\bm{\Sigma}}^{\frac{1}{2}}\bm{\Sigma}^{-\frac{1}{2}}$, there exists a unique $\widetilde{\mathbf{f}}$, characterized by (4.2), such that $(\widetilde{\mathbf{f}},\widetilde{\mathbf{W}},\widetilde{\bm{\mu}},\widetilde{\bm{\Sigma}})$ gives rise to the same observed distributions of $\mathbf{X}_{e}$. At the same time, any $(\widetilde{\mathbf{f}},\widetilde{\mathbf{W}},\widetilde{\bm{\mu}},\widetilde{\bm{\Sigma}})$ for which this holds for all $e\in[M]_{0}$ is of that form. In other words, the mean $\bm{\mu}$ and covariance $\bm{\Sigma}$ of the basal state are unidentifiable, but the mixing function $\mathbf{f}$ and the relative shift matrix $\mathbf{W}\mathbf{L}$ are identifiable up to affine transformation—provided that the observed training perturbation conditions are sufficiently diverse, as formalised by assumption (iii).
Remark 4.3 (Sufficient diversity).
The matrix product $\mathbf{W}\mathbf{L}\in\mathbb{R}^{{d_{Z}}\times M}$ captures the relative effects of the observed perturbations. Specifically, $(\mathbf{W}\mathbf{L})_{je}$ corresponds to the shift in the $j$th latent $Z_{j}$ resulting from $\mathbf{l}_{e}$, relative to a reference condition $\mathbf{l}_{0}$. Moreover, assumption (iii) of Thm. 4.1 implies
$\min\left\{\mathrm{rank}\left(\widetilde{\mathbf{W}}\right),\mathrm{rank}(\mathbf{L})\right\}\geq\mathrm{rank}(\widetilde{\mathbf{W}}\mathbf{L})={d_{Z}}.$ (4.5)
Hence, sufficient diversity requires at least ${d_{Z}}$ elementary perturbations whose associated shift vectors $\mathbf{w}_{k}\in\mathbb{R}^{d_{Z}}$ are linearly independent, and we must observe at least ${d_{Z}}$ perturbation conditions $\mathbf{l}_{e}$ other than $\mathbf{l}_{0}$ such that the relative perturbation vectors $(\mathbf{l}_{e}-\mathbf{l}_{0})\in\mathbb{R}^{K}$ are linearly independent.
Remark 4.4 (Choice of reference).
The choice of reference environment is arbitrary. Here, we choose $e=0$ as reference without loss of generality. Intuitively, if a perturbation is always present (e.g., $l_{e,1}=1$ for all $e$), then its effects cannot be discerned from the basal state. Therefore, only the effects of the relative perturbations $(\mathbf{l}_{e}-\mathbf{l}_{0})$ can be recovered. In practice, we often have access to an unperturbed, purely observational control condition with $\mathbf{l}_{0}=\bm{0}$.
Remark 4.5 (Deterministic vs noisy mixing.).
The mixing function $\mathbf{f}$ in Thm. 4.1 is deterministic, i.e., does not take a separate noise variable ${\bm{\varepsilon}}$ as input, cf. (3.4). In principle, noise can be appended to $\mathbf{z}^{\mathrm{pert}}$ as additional dimensions that are not influenced by perturbations. However, this increases ${d_{Z}}$ and thus makes it harder to satisfy sufficient diversity (see Remark 4.3). Alternatively, the setting of additive noise, $\mathbf{f}(\mathbf{Z})+{\bm{\varepsilon}}$, can be reduced to the noiseless case (Khemakhem et al., 2020).
4.2 From Identifiability to Extrapolation
Since we aim to make distributional predictions for new perturbations $\mathbf{l}_{\mathrm{test}}$, identifiability is only of intermediary interest. The following result highlights the usefulness of the affine identifiability established in Thm. 4.1 for extrapolation. In particular, the result allows us to uniquely predict the observable effects of certain unseen perturbations—specifically, those which can be expressed as linear combinations of the observed perturbations.
Theorem 4.6 (Extrapolation to span of relative perturbations).
Under the same setting and assumptions as in Thm. 4.1, let $\mathbf{l}_{\mathrm{test}}\in\mathbb{R}^{K}$ be an unseen perturbation label such that
$\left(\mathbf{l}_{\mathrm{test}}-\mathbf{l}_{0}\right)\in\mathrm{span}\left(\left\{\mathbf{l}_{e}-\mathbf{l}_{0}\right\}_{e\in[M]}\right).$ (4.6)
Then the effect of $\mathbf{l}_{\mathrm{test}}$ is uniquely identifiable in the sense that
[…]
The proof of Thm. 4.6 is provided in § A.3.
Remark 4.7 (Additive vs linear perturbations).
For our affine identifiability result (Thm. 4.1), it is not necessary that the mean shifts are of the form $\mathbf{W}\mathbf{l}_{e}$. If we replace $\mathbf{W}\mathbf{l}_{e}$ and $\widetilde{\mathbf{W}}\mathbf{l}_{e}$ in (4.1) with arbitrary shift vectors $\mathbf{c}_{e},\widetilde{\mathbf{c}}_{e}\in\mathbb{R}^{{d_{Z}}}$, the same result can be shown to hold with $\mathbf{W}\mathbf{L}$ and $\widetilde{\mathbf{W}}\mathbf{L}$ replaced by $\mathbf{C}$ and $\widetilde{\mathbf{C}}$, defined as the matrices with columns $(\mathbf{c}_{e}-\mathbf{c}_{0})$ and $(\widetilde{\mathbf{c}}_{e}-\widetilde{\mathbf{c}}_{0})$, respectively. That is, the relative shift vectors are identifiable up to affine transformation, regardless of whether they are linear in $\mathbf{l}$. This has implications, e.g., for the CPA model of Lotfollahi et al. (2023) which includes element-wise nonlinear dose-response functions applied to $\mathbf{l}$, see § B.4 for details. However, linearity is leveraged in the proof of our extrapolation result (Thm. 4.6) where it is used in (4.6) to establish a link between $\mathbf{l}_{\mathrm{test}}$ and the training perturbations. Since only $\mathbf{l}_{\mathrm{test}}$ is observed at test time, the above argument thus cannot easily be extended to the extrapolation setting, as this would require establishing a link between $\mathbf{c}_{\mathrm{test}}$ and the training shifts, all of which are unobserved.
Figure 2: Overview of the Perturbation Distribution Autoencoder (PDAE). In a PDAE, the distribution of a target perturbation condition $h$ (purple) is simulated by encoding, perturbing, and (stochastically) decoding data from a source condition $e$ (blue). Dashed arrows indicate model inputs and green boxes model components with trainable parameters. The learning objective (orange, dotted arrows) consists of: a perturbation loss (bottom), measuring, for all pairs of training domains $(e,h)$, the dissimilarity between the true distribution of $\mathbf{X}_{h}$ and its simulated version based on domain $e$; and a distributional reconstruction loss (top), measuring, for each source domain $e$, the dissimilarity between the true conditional distribution of source observations $\mathbf{X}_{e}$ that are mapped to the same encoding $\widehat{\mathbf{Z}}_{e}$ and the corresponding distribution induced by the stochastic decoder. At test time, the target perturbation label $\mathbf{l}_{h}$ can be replaced with an unseen  $\mathbf{l}_{\mathrm{test}}$ to make predictions.
5 Estimation and Prediction: The Perturbation Distribution Autoencoder (PDAE)
To leverage the extrapolation guarantees of Thm. 4.6, we seek to estimate the identifiable parts of the generative process by matching the distributions of the observed data in (2.1) across all perturbation conditions, as prescribed by condition (4.1) in Thm. 4.1. To this end—rather than fitting a probabilistic model by maximising (a lower bound to) the likelihood, as in most existing methods—we take inspiration from the distributional principal autoencoder (Shen & Meinshausen, 2024a) and adapt it for multi-domain perturbation modelling. Our method, the perturbation distribution autoencoder (PDAE), comprises an encoder, a perturbation matrix, and a (stochastic) decoder, which are trained to maximise the similarity between pairs of true and simulated perturbation distributions, see Fig. 2 for an overview. Intuitively, instead of specifying a (Gaussian, as per Thm. 4.1) prior on the basal state from which to sample latents as in (3.1), we use encoded observations from a source domain as perturbed latents and transform them to perturbed latents from a target domain using the learnt perturbation matrix and the assumed additivity as in (3.2).
5.1 Model Components and Outputs
Encoder.
The encoder $\widehat{\mathbf{g}}:\mathbb{R}^{{d_{X}}}\to\mathbb{R}^{{d_{Z}}}$ maps observations $\mathbf{x}$ to the space of perturbation-relevant latents $\mathbf{z}$. Ideally, it should invert the stochastic mixing function in (3.4) in the sense of recovering the perturbed latent state $\mathbf{z}^{\mathrm{pert}}_{e,i}$ (realization of $\mathbf{Z}^{\mathrm{pert}}_{e,i}$ in (3.2)) from observation $\mathbf{x}_{e,i}$ (realization of $\mathbf{X}_{e,i}$). [^2] This is one of the key differences to other methods like CPA (Lotfollahi et al., 2023) and SAMS-VAE (Bereket & Karaletsos, 2023), which seek an encoder that maps to the latent basal state, regardless of the domain $e$. We therefore denote the encoder outputs by
$\widehat{\mathbf{z}}^{\mathrm{pert}}_{e,i}:=\widehat{\mathbf{g}}(\mathbf{x}_{e,i}),$ (5.1)
and refer to them as estimated perturbed latents.
Perturbation Module.
If the encoder recovers the perturbed latents up to affine transformation, the additivity of perturbation effects assumed in (3.2) allows us to map between the latent distributions underlying different perturbation conditions. Specifically, we use a perturbation model parametrised by a perturbation matrix $\widehat{\mathbf{W}}\in\mathbb{R}^{{d_{Z}}\times K}$ to create synthetic perturbed latents from domain $h$ as:
$\widehat{\mathbf{z}}^{\mathrm{pert}}_{e\to h,i}:=\widehat{\mathbf{z}}^{\mathrm{pert}}_{e,i}-\widehat{\mathbf{W}}\mathbf{l}_{e}+\widehat{\mathbf{W}}\mathbf{l}_{h}=\widehat{\mathbf{z}}^{\mathrm{pert}}_{e,i}+\widehat{\mathbf{W}}(\mathbf{l}_{h}-\mathbf{l}_{e}).$ (5.2)
which can be interpreted as undoing the effects of perturbation $\mathbf{l}_{e}$ (i.e., mapping back to the latent basal state) and then simulating perturbation $\mathbf{l}_{h}$. (For $h=e$, this has no effect and $\mathbf{z}^{\mathrm{pert}}_{e\to h,i}=\mathbf{z}^{\mathrm{pert}}_{e,i}$.)
Decoder.
The decoder $\widehat{\mathbf{f}}:\mathbb{R}^{{d_{Z}}}\times\mathbb{R}^{d_{\epsilon}}\to\mathbb{R}^{{d_{X}}}$ maps estimated perturbed latents $\widehat{\mathbf{z}}^{\mathrm{pert}}$ and noise ${\bm{\varepsilon}}\sim\mathbb{Q}_{\bm{\varepsilon}}$ to observations. When evaluated on synthetic perturbed latents from (5.2), we refer to the corresponding (random) outputs
$\widehat{\mathbf{X}}_{e\to h,i}=\widehat{\mathbf{f}}\left(\widehat{\mathbf{z}}^{\mathrm{pert}}_{e\to h,i},\,{\bm{\varepsilon}}\right)\qquad\text{where}\qquad{\bm{\varepsilon}}\sim\mathbb{Q}_{\bm{\varepsilon}}$ (5.3)
as synthetic observations (from domain $h$ based on domain $e$). For $h=e$, i.e., when evaluated on the estimated perturbed latents from (5.1), we refer to the corresponding (random) outputs as reconstructions,
$\widehat{\mathbf{X}}_{e,i}=\widehat{\mathbf{f}}\left(\widehat{\mathbf{z}}^{\mathrm{pert}}_{e,i},\,{\bm{\varepsilon}}\right)\qquad\text{where}\qquad{\bm{\varepsilon}}\sim\mathbb{Q}_{\bm{\varepsilon}}.$ (5.4)
5.2 Induced Distributions
Conditional Distribution Over Reconstructions.
For any given observation $\mathbf{x}$, we can sample reconstructions from the stochastic decoder by fixing its first argument to the encoding $\widehat{\mathbf{g}}(\mathbf{x})$ and sampling noise from $\mathbb{Q}_{\bm{\varepsilon}}$ as in (5.4). Together, the encoder, decoder, and noise distribution thus induce a distribution over reconstructions, formally given by the push-forward $\widehat{\mathbf{f}}\left(\widehat{\mathbf{g}}(\mathbf{x}),\cdot\right)_{\#}\mathbb{Q}_{\bm{\varepsilon}}$. Since this distribution should ideally match that of $\mathbf{X}|\widehat{\mathbf{g}}(\mathbf{X})=\widehat{\mathbf{g}}(\mathbf{x})$ (i.e., the true distribution of observations that share the same encoding), [^3] For example, if $\widehat{\mathbf{g}}$ is injective, the true reconstruction distribution is a point mass on $\mathbf{x}$, whereas if $\widehat{\mathbf{g}}$ is constant, it is the full (unconditional) distribution of $\mathbf{X}$. In this sense, the distribution of reconstructions captures the information about $\mathbf{x}$ that is left unexplained given only its encoding (Shen & Meinshausen, 2024a). we denote it by ^$\widehat{\mathbb{P}}_{\mathbf{X}|\widehat{\mathbf{g}}(\mathbf{X})=\widehat{\mathbf{g}}(\mathbf{x})}$. (Here, we use the notation $\widehat{\mathbb{P}}$ to denote a distribution induced by our model; it does not refer to an empirical distribution.)
Marginal Perturbation Distributions.
Given a distribution $\mathbb{P}_{e}$ and the corresponding perturbation label $\mathbf{l}_{e}$, our model facilitates sampling synthetic observations for another perturbation condition with label $\mathbf{l}_{h}$ via (5.2) and (5.3). We denote the resulting distribution by ^$\widehat{\mathbb{P}}_{e\to h}$, formally defined as the distribution of
$\widehat{\mathbf{f}}\left(\widehat{\mathbf{g}}\left(\mathbf{X}_{e}\right)+\widehat{\mathbf{W}}\left(\mathbf{l}_{h}-\mathbf{l}_{e}\right),{\bm{\varepsilon}}\right)\qquad\text{where}\qquad\mathbf{X}_{e}\sim\mathbb{P}_{e}\qquad\text{and}\qquad{\bm{\varepsilon}}\sim\mathbb{Q}_{\bm{\varepsilon}}.$ (5.5)
Crucially, both $\mathbf{X}_{e}$ and ${\bm{\varepsilon}}$ are random in (5.5). This contrasts with the conditional distribution over reconstructions, for which the first argument of $\widehat{\mathbf{f}}$ is considered fixed and the only stochasticity stems from the noise ${\bm{\varepsilon}}$. Intuitively, when simulating observations from $\mathbb{P}_{h}$ based on $\mathbb{P}_{e}$ via (5.5), the random variable $\widehat{\mathbf{g}}(\mathbf{X}_{e})-\widehat{\mathbf{W}}\mathbf{l}_{e}$ in the first argument of $\widehat{\mathbf{f}}$ plays the role of a random latent basal state from (3.1).
5.3 Population-Level Learning Objective
At the population level (i.e., assuming access to the true distributions $\mathbb{P}_{e}$), we propose the following learning objective:
$\left(\widehat{\mathbf{g}}^{*},\widehat{\mathbf{f}}^{*},\widehat{\mathbf{W}}^{*}\right)\in\operatorname*{arg\,min}_{\widehat{\mathbf{g}},\widehat{\mathbf{f}},\widehat{\mathbf{W}}}\,\mathcal{L}_{\mathrm{pert}}\left(\widehat{\mathbf{g}},\widehat{\mathbf{f}},\widehat{\mathbf{W}};\left\{\left(\mathbb{P}_{e},\mathbf{l}_{e}\right)\right\}_{e\in[M]_{0}}\right)+\lambda\mathcal{L}_{\text{cond-rec}}\left(\widehat{\mathbf{g}},\widehat{\mathbf{f}};\left\{\mathbb{P}_{e}\right\}_{e\in[M]_{0}}\right),$ (5.6)
where $\mathcal{L}_{\mathrm{pert}}$ is a perturbation loss, which depends on the full model and perturbation labels; $\mathcal{L}_{\text{cond-rec}}$ is a (conditional) reconstruction loss, which only depends on the encoder, decoder, and observed distributions and which may be viewed as a regulariser that is helpful in practice; and $\lambda\geq 0$ is a hyperparameter, which trades off the two losses.
Perturbation Loss.
The perturbation loss $\mathcal{L}_{\mathrm{pert}}$ takes the form of the following sum of pairwise distribution losses
$\mathcal{L}_{\mathrm{pert}}\left(\widehat{\mathbf{g}},\widehat{\mathbf{f}},\widehat{\mathbf{W}};\left\{\left(\mathbb{P}_{e},\mathbf{l}_{e}\right)\right\}_{e\in[M]_{0}}\right)$ $:=\sum_{e,h\in[M]_{0}}d\left(\widehat{\mathbb{P}}_{e\to h},\mathbb{P}_{h}\right)$ (5.7)
where ^$\widehat{\mathbb{P}}_{e\to h}$ depends on $(\mathbb{P}_{e},\mathbf{l}_{e},\mathbf{l}_{h})$ and the model parameters via (5.5); and $d$ is a measure of dissimilarity between distributions. Here, we use the negative expected energy score for $d$, i.e.,
$d\left(\widehat{\mathbb{P}}_{e\to h},\mathbb{P}_{h}\right)=-\mathbb{E}_{\mathbf{X}_{h}\sim\mathbb{P}_{h}}\left[\operatorname{ES}_{\beta}\left(\widehat{\mathbb{P}}_{e\to h},\mathbf{X}_{h}\right)\right],$ (5.8)
where $\operatorname{ES}_{\beta}$ denotes the energy-score (Gneiting & Raftery, 2007), defined as
[…]
For $\beta\in(0,2)$, it is a strictly proper scoring rule, meaning that the expected energy score $\mathbb{E}_{\mathbf{X}}[\operatorname{ES}_{\beta}(\mathbb{P},\mathbf{X})]$ is maximised if and only if $\mathbf{X}\sim\mathbb{P}$, see § B.1 for details on probabilistic forecasting and scoring rules. Combined with its computational simplicity, this property makes the negative expected energy score a popular loss function for distributional regression (Shen & Meinshausen, 2024b, a). It also directly implies the following result.
Corollary 5.1.
The objective in (5.7) is minimised if and only if $\mathbb{P}_{h}=\widehat{\mathbb{P}}_{e\to h}$ for all $e,h\in[M]_{0}$.
The sum in (5.7) also includes the case $h=e$. In this case, enforcing $\mathbb{P}_{e}=\widehat{\mathbb{P}}_{e\to e}$ can be viewed as a marginal reconstruction objective and thus as a weaker constraint than that enforced by $\mathcal{L}_{\text{cond-rec}}$ below.
Conditional Reconstruction Loss.
The conditional reconstruction loss $\mathcal{L}_{\text{cond-rec}}$ is given by the following sum of domain-specific DPA losses proposed by Shen & Meinshausen (2024a),
$\mathcal{L}_{\text{cond-rec}}\left(\widehat{\mathbf{g}},\widehat{\mathbf{f}};\left\{\mathbb{P}_{e}\right\}_{e\in[M]_{0}}\right):=\sum_{e\in[M]_{0}}-\mathbb{E}_{\mathbf{X}_{e}\sim\mathbb{P}_{e}}\left[\operatorname{ES}_{\beta}\left(\widehat{\mathbb{P}}_{\mathbf{X}|\widehat{\mathbf{g}}(\mathbf{X})=\widehat{\mathbf{g}}(\mathbf{X}_{e})},\mathbf{X}_{e}\right)\right],$ (5.10)
where each summand on the RHS of (5.10) involves a conditional energy score (i.e., one whose first argument also depends on $\mathbf{X}_{e}$), see § B.3 for further details and discussion.
5.4 Training
In practice, we do not have access to the true distributions $\mathbb{P}_{e}$, but only to empirical distributions based on the available data in (2.1). Therefore, we approximate expectations w.r.t. $\mathbb{P}_{e}$ and $\mathbb{P}_{h}$ in (5.8), (5.9), and (5.10) with Monte Carlo estimates based on mini-batches of (2.1), and use a single draw of ${\bm{\varepsilon}}$ to approximate expectations w.r.t. $\mathbb{Q}_{\bm{\varepsilon}}$. Moreover, for a fixed encoder, the optimal perturbation matrix is available in closed form and given by the ordinary least squares solution to regressing the domain-specific means $\widehat{\bm{\mu}}^{\mathbf{z}}_{e}:=\frac{1}{N_{e}}\sum_{i=1}^{N_{e}}\widehat{\mathbf{z}}^{\mathrm{pert}}_{e,i}$ on the corresponding perturbation labels $\mathbf{l}_{e}$. We, therefore, write $\widehat{\mathbf{W}}$ as a function of $\widehat{\mathbf{g}}$ and optimise (5.6) with respect to the parameters of $\widehat{\mathbf{g}}$ and $\widehat{\mathbf{f}}$ using stochastic gradient descent (Kingma & Ba, 2015). To avoid encouraging the encoder to capture latent dimensions of maximal variation as in the DPA (Shen & Meinshausen, 2024a) rather than focus on the perturbation-relevant latents $\mathbf{Z}$, we do not update the encoder parameters using (the gradient of) the reconstruction loss.
5.5 Prediction
To sample from the predicted distribution for a new perturbation label $\mathbf{l}_{\mathrm{test}}$, we use our model to compute the synthetic perturbed test latents $\widehat{\mathbf{z}}^{\mathrm{pert}}_{e\to\mathrm{test},i}$ for all $e\in[M]_{0}$ and all $i\in[N_{e}]$ via (5.2), and then sample the corresponding synthetic test observations $\widehat{\mathbf{X}}_{e\to\mathrm{test},i}$ according to (5.3). In other words, our estimate of $\mathbb{P}_{\mathrm{test}}$ is the empirical version of the pooled domain-specific marginal perturbation distributions $\mathbb{P}_{e\to\mathrm{test}}$ induced by our model for $\mathbf{l}_{\mathrm{test}}$, i.e.,
^$\widehat{\mathbb{P}}_{\mathrm{test}}=\frac{1}{M+1}\sum_{e\in[M]_{0}}\widehat{\mathbb{P}}_{e\to\mathrm{test}}.$ (5.11)
6 Relation to Prior Work
Perturbation Modelling.
Our model from § 3 is closely inspired by prior work on modelling combinations of genetic or chemical perturbations, particularly the compositional perturbation autoencoder (CPA; Lotfollahi et al., 2019). CPA additionally includes dose-response curves that allow for a nonlinear element-wise rescaling of the perturbation labels $\mathbf{l}$ and employs an adversarial autoencoder for estimation, see § B.4 for further details. Other prior works such as SVAE+ (Lopez et al., 2023) and SAMS-VAE (Bereket & Karaletsos, 2023) instead fit a generative model to the data by maximising the evidence lower bound (ELBO; a lower bound to the marginal likelihood) using variational inference, typically within the variational autoencoder (VAE) framework (Kingma & Welling, 2014; Rezende et al., 2014). Since computing the ELBO requires evaluating the likelihood $p(\mathbf{x}|\mathbf{z})$, this approach generally comes with parametric assumptions. In contrast, our PDAE approach is based on classical (non-variational) autoencoders and is trained using the energy score, which only requires being able to sample from the stochastic decoder. Both SVAE+ and SAMS-VAE additionally impose a sparsity assumption on the effect of perturbations on the latents. SAMS-VAE adopts the same compositional perturbation model as CPA (and the present work) wherein the effect of perturbations can be decomposed into a sum of effects of elementary perturbations, as captured by the linear form $\mathbf{W}\mathbf{l}$. In contrast, SVAE+ extends work on identifiable VAEs with sparsely changing mechanisms (Lachapelle et al., 2022) and models each perturbation condition separately. Importantly, unlike the present work, none of these prior works on perturbation modelling provide theoretical identifiability or extrapolation guarantees.
Relation to Causal Models.
In the field of causal inference, experimental data resulting from perturbations is modelled via interventions in an underlying causal model. For example, in the structural causal model (SCM; Pearl, 2009) framework, interventions modify a set of assignments, which determine each variable from its direct causes and unexplained noise. In general, our model for the effect of perturbations in latent space (§ 3) differs from how interventions are treated in SCMs. However, as detailed in Appx. C, if we restrict our attention to linear SCMs, then the class of shift interventions can be viewed as a special case of our model with $K={d_{Z}}$ elementary perturbations and a particular choice of perturbation matrix $\mathbf{W}$. In this sense, our approach may be interpreted as causal representation learning (Schölkopf et al., 2021) with a linear latent causal model (Squires et al., 2023; Buchholz et al., 2024) and generalized shift interventions (Zhang et al., 2024). In fact, Zhang et al. (2024) also consider the task of extrapolating to unseen combinations of genetic perturbations from a causal representation learning perspective by modelling elementary perturbations as soft interventions (shifts in a nonlinear causal model) on individual latent variables.
Identifiable Representation Learning for Extrapolation.
Saengkyongam et al. (2024, §4) consider an identifiability problem similar to the one studied in § 4 of the present work and also leverage affine identifiability to derive out-of-support extrapolation guarantees. However, their model (and proof strategy) differs from ours, for example, in that they assume the training support of the perturbation labels to contain a non-empty open subset of $\mathbb{R}^{K}$.
7 Experiments
We present preliminary empirical evidence that our approach can outperform existing methods and baselines at distributional perturbation extrapolation. As we focus on a simple, controlled setting with synthetic data, our current results should be viewed as proof of concept, rather than a comprehensive empirical study.
Methods.
We compare our approach with CPA (Lotfollahi et al., 2023) and the following baselines:
  • Pool All:
    pool all training observations;
  • Pseudobulking:
    pool only training data arising from individual perturbations involved in the combination to be predicted (e.g., pool data from $\mathbf{l}_{1}$ and $\mathbf{l}_{2}$ below for predicting $\mathbf{l}_{\mathrm{test}}=(1,1,0)^{\top}$);
  • Linear Regression:
    linearly regress the environment-specific means of observations $\bm{\mu}^{\mathbf{x}}_{e}\in\mathbb{R}^{d_{X}}$ on $\mathbf{l}_{e}$ and use the resulting model to predict the test mean $\bm{\mu}_{\mathrm{test}}^{\mathbf{x}}$ from $\mathbf{l}_{\mathrm{test}}$.
The former two were also used by Lotfollahi et al. (2023); we propose the latter as an additional baseline. Since CPA and linear regression only predict a mean, in these cases, we apply the test mean shift to all observations in the reference condition to simulate the resulting test distribution and use this to also compute distributional similarity metrics.
Metrics.
To assess the distributional fit between predicted and true (empirical) test distributions, we use the energy distance (ED; Székely & Rizzo, 2013), i.e., twice the normalized negative expected (marginal) energy score, and the maximum mean discrepancy (MMD; Gretton et al., 2012) with Gaussian kernel and bandwidth chosen by the median heuristic, see § B.2 for more on measures of distributional similarity. Since some methods are only desgined to predict the mean, we also report the L2 norm of the difference between the predicted and true mean, $\left\lVert\bm{\mu}^{\mathbf{x}}-\widehat{\bm{\mu}}^{\mathbf{x}}\right\rVert$.
Table 1: Results on Synthetic 2D Data. We report average performances over 6 ID and OOD test cases each; shown are the mean $\pm$ one standard deviation over 5 random seeds. Best results highlighted in bold. Theoretical guarantees only apply to ID test cases, i.e., methods cannot be expected to perform well on OOD test cases.
Method In-Distribution (ID) Test Out-of-Distribution (OOD) Test
ED ($\downarrow$) MMD ($\downarrow$) $\left\lVert\bm{\mu}_{\mathbf{x}}-\widehat{\bm{\mu}}^{\mathbf{x}}\right\rVert_{2}$ ($\downarrow$) ED ($\downarrow$) MMD ($\downarrow$) $\left\lVert\bm{\mu}_{\mathbf{x}}-\widehat{\bm{\mu}}^{\mathbf{x}}\right\rVert_{2}$ ($\downarrow$)
Pool All 0.852 $\pm$ .004 1.124 $\pm$ .006 0.385 $\pm$ .001 2.252 $\pm$ .004 3.222 $\pm$ .008 1.488 $\pm$ .002
Pseudobulking 0.723 $\pm$ .003 1.588 $\pm$ .010 0.441 $\pm$ .001 1.619 $\pm$ .005 3.229 $\pm$ .009 1.201 $\pm$ .003
Linear Regression 0.138 $\pm$ .002 0.613 $\pm$ .010 0.215 $\pm$ .002 0.876 $\pm$ .008 2.865 $\pm$ .026 0.770 $\pm$ .005
CPA 0.339 $\pm$ .126 1.411 $\pm$ .363 0.372 $\pm$ .066 1.404 $\pm$ .177 3.706 $\pm$ .343 1.109 $\pm$ .090
PDAE (Ours) 0.001 $\pm$ .000 0.005 $\pm$ .001 0.016 $\pm$ .003 0.113 $\pm$ .020 0.491 $\pm$ .091 0.231 $\pm$ .021
Data.
For ease of visualisation, we first consider the case of ${d_{Z}}\!=\!{d_{X}}\!=\!2$-dimensional latents and observations. The base distribution $\mathbb{P}_{\mathbf{Z}}$ is a zero-mean, isotropic Gaussian with standard deviation $\sigma\!=\!0.25$. We consider $K\!=\!3$ elementary perturbations with associated shift vectors $\mathbf{w}_{1}\!=\!(1,0)^{\top}$, $\mathbf{w}_{2}\!=\!(0,1)^{\top}$, and $\mathbf{w}_{3}\!=\!(1,1)^{\top}$ and create $M\!+\!1\!=4$ training domains with perturbation labels $\mathbf{l}_{0}\!=\!(0,0,0)^{\top}$, $\mathbf{l}_{1}\!=\!(1,0,0)^{\top}$, $\mathbf{l}_{2}\!=\!(0,1,0)^{\top}$, and $\mathbf{l}_{3}\!=\!(0,0,1)^{\top}$. In this case, the sufficient diversity condition (iii) used to establish affine identifiability in Thm. 4.1 is satisfied. For testing, we distinguish between two scenarios: in-distribution (ID) and out-of-distribution (OOD) test cases, depending on whether the support of the distribution of perturbed latents resulting form $\mathbf{l}_{\mathrm{test}}$ lies (mostly) within or outside the support of perturbed training latents (i.e., the inputs “seen” by the decoder during training), respectively. We consider six ID and OOD test cases each, see Appx. D for a list of the exact test labels $\mathbf{l}_{\mathrm{test}}^{\text{ID}}$ and $\mathbf{l}_{\mathrm{test}}^{\text{OOD}}$. To generate observations, we use the complex exponential $\mathbf{x}\!=\!\mathbf{f}(\mathbf{z})\!=\!\mathrm{e}^{z_{1}}(\cos z_{2},\sin z_{2})$ as a deterministic nonlinear mixing function. The resulting training datasets, along with an example of an ID and an OOD test case are in shown in Fig. 3 (left); see Fig. 4 in Appx. D for a complete visualization including all twelve test cases.
Experimental Details.
We generate $N_{e}=2^{14}$ observations for each domain. Both CPA and PDAE use $4$-hidden-layer MLPs with $64$ hidden units as encoders and decoders and are trained for $2000$ epochs using a batch size of $2^{12}$. For PDAE, we set $\beta=1$ in (5.9) and use a learning rate of $0.005$. Since the ground truth mixing function is deterministic, we use a reconstruction loss weight of $\lambda=0$ and no noise dimensions ($\hat{d_{\epsilon}}=0$). For CPA, we select the learning rates of the autoencoder and adversarial classsifier, as well as the latent bottleneck dimension using a grid search based on performance on a held out portion of the training data, as detailed in Appx. D. All other hyperparameters for CPA are set to their default values. We report results for average performance across the six ID and OOD test cases over five repetitions of randomly generating data and initializing, training, and evaluating models.
Figure 3: Qualitative Comparison of PDAE and CPA on Synthetic 2D Data. Rows correspond to latent space (top) and observation space (bottom). Columns show the ground truth data (left), PDAE predictions (center), and CPA predictions (right). Training domains are shown in grey, an in-distribution (ID) test case with $\mathbf{l}^{\textsc{id}}_{\mathrm{test}}\!=\!(1,1,0)^{\top}$ (which overlaps with one of the training domains) in blue, and an out-of-distribution (OOD) test case with $\mathbf{l}_{\mathrm{test}}^{\textsc{ood}}\!=\!(1,0,1)^{\top}$ in orange. All plots show kernel density estimates of the distributions; crosses (x) indicate the corresponding means. As can be seen, PDAE recovers an affine transformation of the true latents (top, center) leading to accurate distributional predictions for the training and ID test domain (bottom, center). However, the OOD test domain is mapped to a part of the latent space not seen during training (top, center). As a result, the corresponding decoder output does not accurately match the true OOD distribution (bottom left vs center). CPA accurately predicts the means of the training distribution (bottom right, black crosses) but does not recover the true latents up to an affine transformation (top right). Moreover, the predicted distributions for CPA do not match the ground truth particularly well, particularly for the test conditions (bottom left vs right). However, recall that—unlike PDAE—CPA is not trained for distributional reconstruction, see § B.4 for details.
Results.
The quantitative results are summarized in Tab. 1; see also Fig. 3 and its caption for qualitative results. For the ID test setting, PDAE performs best, achieving near-perfect distributional fit. CPA outperforms the pooling and pseudobulking baselines (consistent with results reported by Lotfollahi et al. (2023) on gene perturbation data), but does substantially worse than the linear regression and PDAE at both mean and distributional prediction. For the OOD test setting, all methods perform much worse, with PDAE yielding the least bad performance. This failure for the OOD case is expected. As visualised in Fig. 3, despite learning a good representation and perturbation model (as evident from the strong ID test performance), the decoder did not encounter inputs similar to the perturbed test latents during training and has thus not learnt which part of the observation space to map them to.
Noisy Observations.
In Appx. E, we additionally consider a modified version of the aforementioned experiment with ${d_{X}}=10$ (instead of ${d_{X}}=2$) observed dimensions by concatenating ${d_{\epsilon}}=8$ additional dimensions of mean-zero Gaussian noise and investigate the behaviour for increasing noise variance (i.e., decreasing signal-to-noise ratio). In this case, using a non-zero weight ($\lambda>0$) for the conditional reconstruction loss in (5.10) is helpful and we find PDAE to outperform other methods for low to medium noise variances, see Appx. E for details.
8 Discussion
Single-Cell vs Population-Level Effects.
Since cells are typically destroyed during measurement, each unit is only observed for one perturbation condition. As a result, we do not have access to paired data, which would be required to establish ground-truth single-cell level effects. Instead, we can only observe the effects of perturbations at the population level. Thus, we consider distributional perturbation extrapolation (§ 2) a natural formulation of the objective. Our theoretical guarantees and learning objective are formulated accordingly in distributional terms. This contrasts with some prior works, which, despite a lack of ground truth training data, pursue the (seemingly infeasible) task of making counterfactual predictions at the single-cell level. While our model can, in principle, also make such predictions, we stress that they are not falsifiable from empirical data and that our guarantees do not extend to the single-cell resolution. For this reason, we consider distributional perturbation extrapolation a more meaningful and feasible task formulation.
Is Gaussianity Necessary?
In Thm. 4.1, Gaussianity of $\mathbb{P}_{\mathbf{Z}}$ is assumed to prove identifiability. While this is a sufficient condition, it may not be necessary and can possibly be relaxed. Indeed, our method does not explicitly enforce any particular latent distribution, yet still yields strong performance. However, to encourage normality, we could add a regularizer that penalises the KL between the encoded latent basal state distribution and an isotropic Gaussian prior—akin to the KL loss in VAEs. Moreover, preliminary empirical evidence suggests that training on simulated data generated from a Laplacian basal state distribution also yields identifiable latent spaces and perturbation effects.
Decoder Extrapolation and Uncertainty Quantification.
We consider the requirement for the decoder to generalize to new inputs the biggest open problem for distributional perturbation extrapolation. While this issue is absent from our theory, where $\mathbb{P}_{\mathbf{Z}}$ is Gaussian and thus has full support, it can pose major challenges in practice when learning from finite data. As discussed at the end of § 7, the issue of decoder extrapolation is orthogonal to learning the correct representation and perturbation model and thus appears fundamental. If the distribution of perturbed test latents lies fully or partially outside the training support, quantifying the uncertainty associated with different parts of the predicted distribution in observation space would be desirable, e.g., by measuring the overlap in supports or relying on methods for outlier/anomaly detection.
Future Work.
Besides addressing, or at least quantifying, (the lack of) decoder extrapolation, future work should evaluate the proposed approach on more complex, noisy real-world data; extend our theoretical results to partially identifiable settings; and pursue extensions that incorporate covariates and nonlinear dose-response functions.
Appendices
[…]
Appendix A Proofs
A.1 Proof of Thm. 4.1
See 4.1
Proof.
Let $p_{e}$ and $\widetilde{p}_{e}$ denote the densities of
$\mathbf{Z}_{e}:=\mathbf{Z}+\mathbf{W}\mathbf{l}_{e}$ (A.1)
and
$\widetilde{\mathbf{Z}}_{e}:=\widetilde{\mathbf{Z}}+\widetilde{\mathbf{W}}\mathbf{l}_{e},$ (A.2)
respectively. Due to (4.1), $\mathbf{f}$ and $\widetilde{\mathbf{f}}$ have the same image. Thus, by the invertibility assumption (i), the function $\mathbf{h}:=\widetilde{\mathbf{f}}^{-1}\circ\mathbf{f}:\mathbb{R}^{{d_{Z}}}\to\mathbb{R}^{d_{Z}}$ is a well-defined $C^{2}$-diffeomorphism. The change of variable formula applied to
[…]
then yields for all $e$ and all $\mathbf{z}$:
$p_{e}(\mathbf{z})=\widetilde{p}_{e}\left(\mathbf{h}(\mathbf{z})\right)\left|\det\mathbf{J}_{\mathbf{h}}(\mathbf{z})\right|,$ (A.4)
where $\mathbf{J}_{\mathbf{h}}(\mathbf{z})$ denotes the Jacobian of $\mathbf{h}$. By taking logarithms of (A.4) and contrasting domain $e$ with a reference domain with $e=0$, the determinant terms cancel and we obtain for all $e$ and all $\mathbf{z}$:
$\log p_{e}(\mathbf{z})-\log p_{0}(\mathbf{z})=\log\widetilde{p}_{e}\left(\mathbf{h}(\mathbf{z})\right)-\log\widetilde{p}_{0}\left(\mathbf{h}(\mathbf{z})\right).$ (A.5)
Next, denote the densities of $\mathbf{Z}$ and $\widetilde{\mathbf{Z}}$ by $p$ and $\widetilde{p}$, respectively. From (A.1) and (A.2) it then follows that, for all $e$ and all $\mathbf{z}$, $p_{e}$ and $\widetilde{p}_{e}$ can respectively be expressed in terms of $p$ and $\widetilde{p}$ as follows:
$p_{e}(\mathbf{z})$ $=p\left(\mathbf{z}-\mathbf{W}\mathbf{l}_{e}\right),$ (A.6)
$\widetilde{p}_{e}(\mathbf{z})$ $=\widetilde{p}\left(\mathbf{z}-\widetilde{\mathbf{W}}\mathbf{l}_{e}\right).$ (A.7)
By substituting these expressions in (A.5), we obtain for all $e$ and all $\mathbf{z}$:
$\log p(\mathbf{z}-\mathbf{W}\mathbf{l}_{e})-\log p(\mathbf{z}-\mathbf{W}\mathbf{l}_{0})=\log\widetilde{p}\left(\mathbf{h}(\mathbf{z})-\widetilde{\mathbf{W}}\mathbf{l}_{e}\right)-\log\widetilde{p}\left(\mathbf{h}(\mathbf{z})-\widetilde{\mathbf{W}}\mathbf{l}_{0}\right).$ (A.8)
By the Gaussianity assumption (ii), the contrast of log-densities in (A.8) takes the following form for all $e$ and all $\mathbf{z}$:
$(\mathbf{l}_{e}-\mathbf{l}_{0})^{\top}\mathbf{W}^{\top}\bm{\Sigma}^{-1}(\mathbf{z}-\bm{\mu})$ $-\frac{1}{2}(\mathbf{l}_{e}-\mathbf{l}_{0})^{\top}\mathbf{W}^{\top}\bm{\Sigma}^{-1}\mathbf{W}(\mathbf{l}_{e}+\mathbf{l}_{0})$ (A.9)
=\,= $(\mathbf{l}_{e}-\mathbf{l}_{0})^{\top}\widetilde{\mathbf{W}}^{\top}\widetilde{\bm{\Sigma}}^{-1}\left(\mathbf{h}(\mathbf{z})-\widetilde{\bm{\mu}}\right)$ $-\frac{1}{2}(\mathbf{l}_{e}-\mathbf{l}_{0})^{\top}\widetilde{\mathbf{W}}^{\top}\widetilde{\bm{\Sigma}}^{-1}\widetilde{\mathbf{W}}(\mathbf{l}_{e}+\mathbf{l}_{0})$ (A.10)
We aim to show that the two representations are related by an affine transformation, i.e., that all second-order derivatives of $\mathbf{h}$ are zero everywhere. Taking gradients w.r.t. $\mathbf{z}$ yields for all $e$ and all $\mathbf{z}$:
$(\mathbf{l}_{e}-\mathbf{l}_{0})^{\top}\mathbf{W}^{\top}\bm{\Sigma}^{-1}=(\mathbf{l}_{e}-\mathbf{l}_{0})^{\top}\widetilde{\mathbf{W}}^{\top}\widetilde{\bm{\Sigma}}^{-1}\mathbf{J}_{\mathbf{h}}(\mathbf{z}).$ (A.11)
Let $\mathbf{L}\in\mathbb{R}^{K\times M}$ be the matrix with columns $\mathbf{l}_{e}-\mathbf{l}_{0}$ for $e\in[M]$. Then, for all $\mathbf{z}$:
$\mathbf{L}^{\top}\mathbf{W}^{\top}\bm{\Sigma}^{-1}=\mathbf{L}^{\top}\widetilde{\mathbf{W}}^{\top}\widetilde{\bm{\Sigma}}^{-1}\mathbf{J}_{\mathbf{h}}(\mathbf{z}).$ (A.12)
Differentiating once more w.r.t. $\mathbf{z}$ yields for all $\mathbf{z}$:
$\bm{0}=\mathbf{L}^{\top}\widetilde{\mathbf{W}}^{\top}\widetilde{\bm{\Sigma}}^{-1}\mathbf{H}_{\mathbf{h}}(\mathbf{z}),$ (A.13)
where the 3-tensor $\mathbf{H}_{\mathbf{h}}(\mathbf{z})\in\mathbb{R}^{{d_{Z}}\times{d_{Z}}\times{d_{Z}}}$ denotes the Hessian of $\mathbf{h}$, i.e., for all $i,j\in[{d_{Z}}]$ and all $\mathbf{z}$
$\bm{0}=\mathbf{L}^{\top}\widetilde{\mathbf{W}}^{\top}\widetilde{\bm{\Sigma}}^{-1}\frac{\partial^{2}}{\partial z_{i}\partial z_{j}}\mathbf{h}(\mathbf{z}).$ (A.14)
By assumption (iii), the matrix $\mathbf{L}^{\top}\widetilde{\mathbf{W}}^{\top}$ has full column rank and thus a left inverse, i.e., there exists $\mathbf{V}\in\mathbb{R}^{{d_{Z}}\times M}$ such that $\mathbf{V}\mathbf{L}^{\top}\widetilde{\mathbf{W}}^{\top}=\mathbf{I}_{{d_{Z}}}$. Multiplying (A.14) on the left by $\widetilde{\bm{\Sigma}}\mathbf{V}$ then yields for all $i,j\in[{d_{Z}}]$ and all $\mathbf{z}$:
$\frac{\partial^{2}}{\partial z_{i}\partial z_{j}}\mathbf{h}(\mathbf{z})=\bm{0}.$ (A.15)
This implies that $\mathbf{h}$ must be affine, i.e., there exist $\mathbf{A}\in\mathbb{R}^{{d_{Z}}\times{d_{Z}}}$ and $\mathbf{b}\in\mathbb{R}^{d_{Z}}$ such that for all $\mathbf{z}$:
$\mathbf{h}(\mathbf{z})=\mathbf{A}\mathbf{z}+\mathbf{b}.$ (A.16)
Further, since $\mathbf{h}$ is invertible, $\mathbf{A}$ must be invertible.
[…]
$\mathcal{N}\left(\widetilde{\bm{\mu}}+\widetilde{\mathbf{W}}\mathbf{l}_{e},\widetilde{\bm{\Sigma}}\right)=\mathcal{N}\left(\mathbf{A}(\bm{\mu}+\mathbf{W}\mathbf{l}_{e})+\mathbf{b},\mathbf{A}\bm{\Sigma}\mathbf{A}^{\top}\right).$ (A.17)
By equating the covariances, we find
$\mathbf{A}\bm{\Sigma}\mathbf{A}^{\top}=\widetilde{\bm{\Sigma}}\qquad\implies\qquad\mathbf{A}=\widetilde{\bm{\Sigma}}^{\frac{1}{2}}\bm{\Sigma}^{-\frac{1}{2}},$ (A.18)
where the matrix square roots exist and are unique since $\bm{\Sigma}$ and $\widetilde{\bm{\Sigma}}$ are symmetric and positive definite. By equating the means, we obtain for all $e$:
$\widetilde{\bm{\mu}}+\widetilde{\mathbf{W}}\mathbf{l}_{e}=\mathbf{A}(\bm{\mu}+\mathbf{W}\mathbf{l}_{e})+\mathbf{b}\qquad\iff\qquad(\widetilde{\mathbf{W}}-\mathbf{A}\mathbf{W})\mathbf{l}_{e}=\mathbf{A}\bm{\mu}+\mathbf{b}-\widetilde{\bm{\mu}}.$ (A.19)
By contrasting (A.19) for all $e\in[M]$ with $e=0$ as before, we obtain
$(\widetilde{\mathbf{W}}-\mathbf{A}\mathbf{W})\mathbf{L}=\bm{0}\qquad\iff\qquad\widetilde{\mathbf{W}}\mathbf{L}=\mathbf{A}\mathbf{W}\mathbf{L}.$ (A.20)
Finally, by choosing $e=0$ in (A.19) we obtain the desired expression for $\mathbf{b}$,
$\mathbf{b}=\widetilde{\bm{\mu}}-\mathbf{A}\bm{\mu}+(\widetilde{\mathbf{W}}-\mathbf{A}\mathbf{W})\mathbf{l}_{0}.$ (A.21)
This completes the proof. ∎
A.2 Proof of Cor. 4.2
See 4.2
Proof.
If $\mathrm{rank}(\mathbf{L})=K$, then $\mathbf{L}$ has a right inverse, i.e., there exists $\mathbf{K}\in\mathbb{R}^{M\times K}$ such that $\mathbf{L}\mathbf{K}=\mathbf{I}_{K}$. Right multiplication of (A.20) by $\mathbf{K}$ then yields
$\widetilde{\mathbf{W}}=\mathbf{A}\mathbf{W}$ (A.22)
Finally, substitution of (A.22) into (A.21) yields
$\mathbf{b}=\widetilde{\bm{\mu}}-\mathbf{A}\bm{\mu}.$ (A.23)
A.3 Proof of Thm. 4.6
See 4.6
Proof.
With $\mathbf{h}=\widetilde{\mathbf{f}}^{-1}\circ\mathbf{f}$, the condition in (4.7) is equivalent to
[…]
By Thm. 4.1, we have $\mathbf{h}(\mathbf{z})=\mathbf{A}\mathbf{z}+\mathbf{b}$ for
$\mathbf{A}$ $=\widetilde{\bm{\Sigma}}^{\frac{1}{2}}\bm{\Sigma}^{-\frac{1}{2}},$ (A.25)
$\mathbf{b}$ $=\widetilde{\bm{\mu}}-\mathbf{A}\bm{\mu}+(\widetilde{\mathbf{W}}-\mathbf{A}\mathbf{W})\mathbf{l}_{0}.$ (A.26)
Together with $\mathbf{Z}\sim\mathcal{N}(\bm{\mu},\bm{\Sigma})$, this lets us compute the distribution of the LHS of (A.24) as:
$\mathbf{h}(\mathbf{Z}+\mathbf{W}\mathbf{l}_{\mathrm{test}})=\mathbf{A}(\mathbf{Z}+\mathbf{W}\mathbf{l}_{\mathrm{test}})+\mathbf{b}\sim\mathcal{N}\left(\mathbf{A}\bm{\mu}+\mathbf{A}\mathbf{W}\mathbf{l}_{\mathrm{test}}+\mathbf{b},\mathbf{A}\bm{\Sigma}\mathbf{A}^{\top}\right).$ (A.27)
Similarly, since $\widetilde{\mathbf{Z}}\sim\mathcal{N}\left(\widetilde{\bm{\mu}},\widetilde{\bm{\Sigma}}\right)$, the distribution of the RHS of (A.24) is given by
$\widetilde{\mathbf{Z}}+\widetilde{\mathbf{W}}\mathbf{l}_{\mathrm{test}}\sim\mathcal{N}\left(\widetilde{\bm{\mu}}+\widetilde{\mathbf{W}}\mathbf{l}_{\mathrm{test}},\widetilde{\bm{\Sigma}}\right).$ (A.28)
From (A.25), it follows directly that
$\mathbf{A}\bm{\Sigma}\mathbf{A}^{\top}=\widetilde{\bm{\Sigma}}.$ (A.29)
To complete the proof, it thus remains to show that the means of (A.27) and (A.28) are equal, i.e.,
$\mathbf{A}\bm{\mu}+\mathbf{A}\mathbf{W}\mathbf{l}_{\mathrm{test}}+\mathbf{b}=\widetilde{\bm{\mu}}+\widetilde{\mathbf{W}}\mathbf{l}_{\mathrm{test}}.$ (A.30)
Starting from the LHS of (A.30), by substituting the expression for $\mathbf{b}$ from (A.26) we obtain:
$\mathbf{A}\bm{\mu}+\mathbf{A}\mathbf{W}\mathbf{l}_{\mathrm{test}}+\mathbf{b}$ $=\mathbf{A}\bm{\mu}+\mathbf{A}\mathbf{W}\mathbf{l}_{\mathrm{test}}+\widetilde{\bm{\mu}}-\mathbf{A}\bm{\mu}+(\widetilde{\mathbf{W}}-\mathbf{A}\mathbf{W})\mathbf{l}_{0}$ (A.31)
$=\widetilde{\bm{\mu}}+\mathbf{A}\mathbf{W}(\mathbf{l}_{\mathrm{test}}-\mathbf{l}_{0})+\widetilde{\mathbf{W}}\mathbf{l}_{0}$ (A.32)
Next, by (4.6), i.e., the assumption that $(\mathbf{l}_{\mathrm{test}}-\mathbf{l}_{0})$ lies in the span of $\{\mathbf{l}_{e}-\mathbf{l}_{0}\}_{e\in[M]}$, there exists $\bm{\alpha}\in\mathbb{R}^{M}$ such that
$\mathbf{l}_{\mathrm{test}}-\mathbf{l}_{0}=\sum_{e\in[M]}\alpha_{e}(\mathbf{l}_{e}-\mathbf{l}_{0})=\mathbf{L}\bm{\alpha}$ (A.33)
where, as before, $\mathbf{L}\in\mathbb{R}^{K\times M}$ is the matrix with columns $(\mathbf{l}_{e}-\mathbf{l}_{0})$. Substituting (A.33) into (A.32) yields
$\mathbf{A}\bm{\mu}+\mathbf{A}\mathbf{W}\mathbf{l}_{\mathrm{test}}+\mathbf{b}=\widetilde{\bm{\mu}}+\mathbf{A}\mathbf{W}\mathbf{L}\bm{\alpha}+\widetilde{\mathbf{W}}\mathbf{l}_{0}.$ (A.34)
Finally, it follows from Thm. 4.1 that
$\mathbf{A}\mathbf{W}\mathbf{L}=\widetilde{\mathbf{W}}\mathbf{L}.$ (A.35)
which upon substituion into (A.34) yields the desired equality from (A.30)
$\mathbf{A}\bm{\mu}+\mathbf{A}\mathbf{W}\mathbf{l}_{\mathrm{test}}+\mathbf{b}$ $=\widetilde{\bm{\mu}}+\widetilde{\mathbf{W}}\mathbf{L}\bm{\alpha}+\widetilde{\mathbf{W}}\mathbf{l}_{0}$ (A.36)
$=\widetilde{\bm{\mu}}+\widetilde{\mathbf{W}}(\mathbf{L}\bm{\alpha}+\mathbf{l}_{0})$ (A.37)
$=\widetilde{\bm{\mu}}+\widetilde{\mathbf{W}}\mathbf{l}_{\mathrm{test}}.$ (A.38)
This completes the proof. ∎
Appendix B Additional Background Material and Related Work
Since the problem of interest (§ 2) involves making distributional predictions for new perturbation conditions, we review some basics of probabilistic forecasting (§ B.1) and measuring the similarity between two distributions (§ B.2), in our case typically between an empirical distribution and its predicted counterpart. We then turn to representation learning with encoder-decoder architectures (§ B.3), in particular a recent approach that also targets distributional reconstruction. Finally, we cover some prior efforts on perturbation modelling and extrapolation (§ B.4), which we draw inspiration from.
B.1 Probabilistic Forecasting and Scoring Rules
Let $\Omega$ be a sample space, $\mathcal{A}$ a $\sigma$-algebra of subsets of $\Omega$, and $\mathcal{P}$ a convex class of probability measures on $(\Omega,\mathcal{A})$. A probabilistic prediction or probabilistic forecast is a mapping into $\mathcal{P}$, which outputs predictive distributions $\mathbb{P}$ over over outcomes ${x\in\Omega}$. Probabilistic forecasting can thus be viewed as a distributional generalization point prediction (i.e., deterministic forecasting), which maps directly into $\Omega$.
To evaluate, compare, or rank different forecasts, it is useful to assign them a numerical score reflecting their quality. A scoring rule is a function $S:\mathcal{P}\times\Omega\to\mathbb{R}$ that assigns a score $S(\mathbb{P},x)$ to forecast $\mathbb{P}$ if event $x$ materializes, with higher scores corresponding to better forecasts—akin to (negative) loss or cost functions for point predictions. If $x$ is distributed according to $\mathbb{Q}$, we denote the expected score by $S(\mathbb{P},\mathbb{Q})=\mathbb{E}_{x\sim\mathbb{Q}}[S(\mathbb{P},x)]$. The scoring rule is called proper if $S(\mathbb{Q},\mathbb{Q})\geq S(\mathbb{P},\mathbb{Q})$ for all $\mathbb{P}$, and strictly proper if equality holds if and only if $\mathbb{P}=\mathbb{Q}$.
CRPS.
For continuous scalar random variables (i.e., $\Omega=\mathbb{R}$), a popular scoring rule is the continuous ranked probability score (CRPS; Matheson & Winkler, 1976). When $\mathcal{P}$ is the space of Borel probability measures on $\mathbb{R}$ with finite first moment, it is strictly proper and given by: [^4] The original definition by Matheson & Winkler (1976) is $\operatorname{CRPS}(F,x)=-\int_{-\infty}^{\infty}(F(y)-1\{y\geq x\})^{2}\mathrm{d}y$ where $F$ is the CDF of $\mathbb{P}$, but the simpler form in (B.1) has been shown to be equivalent (Baringhaus & Franz, 2004, Lemma 2.2).
[…]
If $\mathbb{P}$ is a point mass, the negative CRPS reduces to the absolute error loss function; it can thus be viewed as a generalization thereof to probabilistic forecasts.
Energy Score.
Gneiting & Raftery (2007) propose the energy score as a multi-variate generalization of the CRPS for vector-valued $\mathbf{x}\in\Omega=\mathbb{R}^{{d_{X}}}$. For $\beta\in(0,2)$, it is defined by
[…]
where $\left\lVert\,\cdot\,\right\rVert_{2}$ denotes the Euclidean (L2) norm. $\operatorname{ES}_{\beta}$ is strictly proper w.r.t. $\mathcal{P}_{\beta}$, the set of Borel probability measures $\mathbb{P}$ for which $\mathbb{E}_{\mathbf{X}\sim\mathbb{P}}\left\lVert\mathbf{X}\right\rVert_{2}^{\beta}$ is finite (Gneiting & Raftery, 2007).
B.2 Assessing Distributional Similarity
Energy Distance.
The expected energy score $\operatorname{ES}_{\beta}(\mathbb{P},\mathbb{Q})=\mathbb{E}_{\mathbf{Y}\sim\mathbb{Q}}[\operatorname{ES}_{\beta}(\mathbb{P},\mathbf{Y})]$ consitutes a measure of similarity between $\mathbb{P}$ and $\mathbb{Q}$ and is closely linked to the energy distance (Székely & Rizzo, 2013):
$\operatorname{ED}_{\beta}(\mathbb{P},\mathbb{Q})$ $=2\mathbb{E}_{\mathbf{X}\sim\mathbb{P},\mathbf{Y}\sim\mathbb{Q}}\left\lVert\mathbf{X}-\mathbf{Y}\right\rVert_{2}^{\beta}-\mathbb{E}_{\mathbf{X},\mathbf{X}^{\prime}\overset{\mathrm{iid}}{\sim}\mathbb{P}}\left\lVert\mathbf{X}-\mathbf{X}^{\prime}\right\rVert_{2}^{\beta}-\mathbb{E}_{\mathbf{Y},\mathbf{Y}^{\prime}\overset{\mathrm{iid}}{\sim}\mathbb{Q}}\left\lVert\mathbf{Y}-\mathbf{Y}^{\prime}\right\rVert_{2}^{\beta}$ (B.3)
$=2\left(\operatorname{ES}_{\beta}(\mathbb{Q},\mathbb{Q})-\operatorname{ES}_{\beta}(\mathbb{P},\mathbb{Q})\right)\geq 0$ (B.4)
with equality if and only if $\mathbb{P}=\mathbb{Q}$, since $\operatorname{ES}_{\beta}$ is a strictly proper scoring rule.
Maximum Mean Discrepancy (MMD).
Another well-known distance between probability measures that is rooted in kernel methods (Schölkopf & Smola, 2002) is the maximum mean discrepancy (MMD; Gretton et al., 2012), which for a positive definite kernel $k:\Omega\times\Omega\to\mathbb{R}$ is given by
$\operatorname{MMD}^{2}_{k}(\mathbb{P},\mathbb{Q})=\mathbb{E}_{X,X^{\prime}\overset{\mathrm{iid}}{\sim}\mathbb{P}}[k(X,X^{\prime})]-2\mathbb{E}_{X\sim\mathbb{P},Y\sim\mathbb{Q}}[k(X,Y)]+\mathbb{E}_{Y,Y^{\prime}\overset{\mathrm{iid}}{\sim}\mathbb{Q}}[k(Y,Y^{\prime})]\,.$ (B.5)
Energy Distance as a Special Case of MMD.
As shown by Sejdinovic et al. (2013), if $k$ in (B.5) is chosen to be the positive-definite distance kernel555induced by the negative-definite semi-metric $\rho(\mathbf{X},\mathbf{Y})=\left\lVert\mathbf{X}-\mathbf{Y}\right\rVert_{2}^{\beta}$ on $\Omega=\mathbb{R}^{{d_{X}}}$ (centered at the origin),
$k_{\beta}(\mathbf{X},\mathbf{Y})=\frac{1}{2}\left(\left\lVert\mathbf{X}\right\rVert_{2}^{\beta}+\left\lVert\mathbf{Y}\right\rVert_{2}^{\beta}-\left\lVert\mathbf{X}-\mathbf{Y}\right\rVert_{2}^{\beta}\right)$ (B.6)
then the energy distance is recovered as a special case of MMD,
$\operatorname{ED}_{\beta}(\mathbb{P},\mathbb{Q})=2\operatorname{MMD}^{2}_{k_{\beta}}(\mathbb{P},\mathbb{Q})\,.$ (B.7)
B.3 Representation Learning
Many modern data sources of interest contain high-dimensional and unstructured observations $\mathbf{x}$, such as audio, video, images, or text. Representation learning aims to transform such data into a more compact, lower-dimensional set of features $\mathbf{z}$ (the representation) which preserves most of the relevant information while making it more easily accessible, e.g., for use in downstream tasks (Bengio et al., 2013). [^6] Representation learning is thus closely related to the classical task of dimension reduction. The former usually refers to nonlinear settings, involves some form of machine learning, and tends to be more focused on usefulness in terms of downstream tasks, rather than, say, explained variance. For example, a representation of images of multi-object scenes could be a list of the contained objects, along with their size, position, colour, etc. A key assumption underlying this endeavour is the so-called manifold hypothesis which posits that high-dimensional natural data tends to lie near a low-dimensional manifold embedded in the high-dimensional ambient space; this idea is also at the heart of several (nonlinear) dimension reduction techniques (Cayton, 2005; Tenenbaum et al., 2000; Saul & Roweis, 2003; Belkin & Niyogi, 2001).
Autoencoder (AE).
An autoencoder (AE; Hinton & Salakhutdinov, 2006; Rumelhart et al., 1986) is a pair of functions $(\mathbf{g},\mathbf{f})$, consisting of an encoder $\mathbf{g}:\mathbb{R}^{{d_{X}}}\to\mathbb{R}^{{d_{Z}}}$ mapping observations $\mathbf{X}\sim\mathbb{P}$ to their representation $\mathbf{Z}:=\mathbf{g}(\mathbf{X})$, and a decoder $\mathbf{f}:\mathbb{R}^{{d_{Z}}}\to\mathbb{R}^{{d_{X}}}$ mapping representations $\mathbf{Z}$ to their reconstructions in observation space, $\widehat{\mathbf{X}}:=\mathbf{f}(\mathbf{Z})=\mathbf{f}(\mathbf{g}(\mathbf{X}))$. Typically, ${d_{Z}}<{d_{X}}$, such that there is a bottleneck and perfect reconstruction is not feasible. Autoencoders are usually trained with a (point-wise, mean) reconstruction objective, i.e., the aim is to minimise the mean squared error
$\mathcal{L}_{\textsc{ae}}(\mathbf{g},\mathbf{f};\mathbb{P}):=\mathbb{E}_{\mathbf{X}\sim\mathbb{P}}\left\lVert\mathbf{X}-\widehat{\mathbf{X}}\right\rVert_{2}^{2}=\mathbb{E}_{\mathbf{X}\sim\mathbb{P}}\left\lVert\mathbf{X}-\mathbf{f}(\mathbf{g}(\mathbf{X}))\right\rVert_{2}^{2}\,,$ (B.8)
w.r.t. both $\mathbf{g}$ and $\mathbf{f}$. As a result, the optimal decoder $\mathbf{f}^{*}_{\textsc{ae}}$ for any fixed encoder choice $\mathbf{g}$ is given by the conditional mean
$\mathbf{f}^{*}_{\textsc{ae}}(\mathbf{z};\mathbf{g},\mathbb{P})=\mathbb{E}_{\mathbf{X}\sim\mathbb{P}}[\mathbf{X}|\mathbf{g}(\mathbf{X})=\mathbf{z}]\,.$ (B.9)
Distributional Principal Autoencoder (DPA).
Since the objective of a standard AE is mean reconstruction, the distribution of reconstructions $\widehat{\mathbf{X}}$ is typically not the same as the distribution of $\mathbf{X}$ (unless the encoder is invertible, i.e, the compression is lossless and perfect reconstruction is feasible, which is usually not the case in practice). To address this, Shen & Meinshausen (2024a) proposed the distributional principal autoencoder (DPA) which targets distributional (rather than mean) reconstruction. A DPA also consists of an encoder-decoder pair $(\mathbf{g},\mathbf{f})$. However, unlike in a standard AE, the DPA decoder $\mathbf{f}:\mathbb{R}^{{d_{Z}}}\times\mathbb{R}^{d_{\epsilon}}\to\mathbb{R}^{{d_{X}}}$ is stochastic and receives an additional noise term ${\bm{\varepsilon}}$ as input, which is sampled from a fixed distribution $\mathbb{Q}_{{\bm{\varepsilon}}}$ such as a standard isotropic Gaussian. The DPA loss function is constructed such that for a fixed encoder $\mathbf{g}$, the optimal DPA decoder $\mathbf{f}^{*}_{\textsc{dpa}}$ maps a given latent embedding $\mathbf{z}$, to the distribution of $\mathbf{X}$, given $\mathbf{g}(\mathbf{X})=\mathbf{z}$, i.e.,
[…]
$\mathcal{L}_{\textsc{dpa}}(\mathbf{g},\mathbf{f};\mathbb{P})$ $=\mathbb{E}_{\mathbf{X}\sim\mathbb{P},{\bm{\varepsilon}}\sim\mathbb{Q}_{{\bm{\varepsilon}}}}\left\lVert\mathbf{X}-\mathbf{f}(\mathbf{g}(\mathbf{X}),{\bm{\varepsilon}})\right\rVert_{2}^{\beta}-\frac{1}{2}\mathbb{E}_{\mathbf{X}\sim\mathbb{P},{\bm{\varepsilon}},{\bm{\varepsilon}}^{\prime}\overset{\mathrm{iid}}{\sim}\mathbb{Q}_{{\bm{\varepsilon}}}}\left\lVert\mathbf{f}(\mathbf{g}(\mathbf{X}),{\bm{\varepsilon}})-\mathbf{f}(\mathbf{g}(\mathbf{X}),{\bm{\varepsilon}}^{\prime})\right\rVert_{2}^{\beta}$ (B.11)
$=-\mathbb{E}_{\mathbf{X}\sim\mathbb{P}}\left[\operatorname{ES}_{\beta}\left(\mathbf{f}\left(\mathbf{g}(\mathbf{X}),\,\cdot\,\right)_{\#}\mathbb{Q}_{{\bm{\varepsilon}}},\mathbf{X}\right)\right]\,,$
where $\mathbf{f}\left(\mathbf{z},\,\cdot\,\right)_{\#}\mathbb{Q}_{{\bm{\varepsilon}}}$ denotes the pushforward distribution of $\mathbb{Q}_{{\bm{\varepsilon}}}$ through the function $\mathbf{f}(\mathbf{z},\,\cdot\,)$, i.e., the distribution of $\mathbf{f}(\mathbf{z},{\bm{\varepsilon}})$ for a fixed $\mathbf{z}$ when ${\bm{\varepsilon}}\sim\mathbb{Q}_{\bm{\varepsilon}}$. In other words, a DPA minimizes the negative expected energy score between $\mathbf{X}$ and the corresponding (stochastic) decoder output, conditional on the encoding of $\mathbf{X}$. Due to this conditioning, the DPA objective differs from an energy distance by a “normalization constant” which depends on the encoder and encourages capturing principal (i.e., variation-minimizing) components, rather than random latent dimensions.
B.4 Perturbation Modelling
Compositional Perturbation Autoencoder (CPA).
Lotfollahi et al. (2023) propose the compositional perturbation autoencoder (CPA) as a model for compositional extrapolation of perturbation data. Specifically, they assume the following model:
$\mathbf{z}^{\mathrm{pert}}$ $=\mathbf{z}^{\mathrm{base}}+\mathbf{W}^{\mathrm{pert}}\begin{pmatrix}h_{1}(l_{1})\\ \dots\\ h_{K}(l_{K})\end{pmatrix}+\sum_{j=1}^{J}\mathbf{W}_{j}^{\text{cov}}\mathbf{c}_{j}\,,$ (B.12)
where $\mathbf{z}^{\mathrm{base}}\sim\mathbb{P}_{\mathbf{Z}}$ denotes an unperturbed basal state; the matrix $\mathbf{W}^{\mathrm{pert}}\in\mathbb{R}^{{d_{Z}}\times K}$ encodes the additive effect of each elementary perturbation; $\{h_{k}:\mathbb{R}\to\mathbb{R}\}_{k\in[K]}$ are unknown, possibly nonlinear dose-response curves; $\{\mathbf{c}_{j}\in\mathbb{R}^{K_{j}}\}_{j\in[J]}$ are observed one-hot vectors capturing $J$ additional discrete covariates, such as cell-types or species; and the matrices $\{\mathbf{W}_{j}^{\text{cov}}\in\mathbb{R}^{{d_{Z}}\times K_{j}}\}_{j\in[J]}$ encode additive covariate-specific effects. Further, the basal state $\mathbf{z}^{\mathrm{base}}$ is assumed to be independent of the perturbation labels $\mathbf{l}$ and covariates $\mathbf{C}=(\mathbf{c}_{1},\dots,\mathbf{c}_{J})$. Observations $\mathbf{x}$ are then drawn from a Gaussian whose mean and variance are determined by the perturbed latent state $\mathbf{z}^{\mathrm{pert}}$,
$\mathbf{x}\sim\mathcal{N}\left(\bm{\mu}\left(\mathbf{z}^{\mathrm{pert}}\right),\sigma^{2}\left(\mathbf{z}^{\mathrm{pert}}\right)\mathbf{I}\right)\,.$ (B.13)
To fit this model, Lotfollahi et al. (2023) employ an adversarial autoencoder (Lample et al., 2017). First, an encoder $\mathbf{g}$ estimates the basal state
$\widehat{\mathbf{z}}^{\mathrm{base}}=\mathbf{g}(\mathbf{x})\,.$ (B.14)
The estimated perturbed latent state $\widehat{\mathbf{z}}^{\mathrm{pert}}$ is then computed according to (B.12) using (B.14) and learnt estimates $\widehat{\mathbf{W}}^{\mathrm{pert}}$, $\{\widehat{h}_{k}\}$, and $\{\widehat{\mathbf{W}}_{l}^{\text{cov}}\}$. Finally, a (deterministic) decoder $\mathbf{f}$ uses $\widehat{\mathbf{z}}^{\mathrm{pert}}$ to compute estimates of the mean and variance in (B.13), i.e., $(\widehat{\bm{\mu}},\widehat{\sigma}^{2})=\mathbf{f}(\widehat{\mathbf{z}}^{\mathrm{pert}})$. All learnable components of the model are trained by minimizing the (Gaussian) negative log-likelihood of the observed data $\mathcal{D}=\{(\mathbf{x}_{i},\mathbf{l}_{i},\mathbf{C}_{i})\}_{i\in[N]}$. To encourage the postulated independence of $\widehat{\mathbf{z}}^{\mathrm{base}}$ and $(\mathbf{l},\mathbf{C})$, an additional adversarial loss is used, which minimizes the predictability of the latter from the former.
Appendix C Relation to Causal Models
In the field of causal inference, experimental data resulting from perturbations is typically modelled via interventions in an underlying causal model. In the structural causal model (SCM) framework of Pearl (2009), interventions modify a set of assignments, which determine each variable as a function of its direct causes and unexplained noise.
Definition C.1 (Acyclic SCM).
An acyclic structural causal model (SCM) $\mathcal{M}=\left({\mathcal{S}},\mathbb{P}_{\mathbf{U}}\right)$ with endogenous variables $\mathbf{V}=\{V_{1},...,V_{n}\}$ and exogenous variables $\mathbf{U}=\{U_{1},...,U_{m}\}$ consists of:
  1. (i)
    a set of structural equations
    ${\mathcal{S}}=\left\{V_{i}:=f_{i}\left(\mathbf{V}_{\mathrm{pa}(i)},\mathbf{U}_{i}\right)\right\}_{i=1}^{n},$ (C.1)
    where $f_{i}$ are deterministic functions; $\mathbf{U}_{i}\subseteq\mathbf{U}$; and $\mathbf{V}_{\mathrm{pa}(i)}\subseteq\mathbf{V}\setminus\{V_{i}\}$ is the set of causal parents of $V_{i}$, such that the induced causal graph with vertices $[n]$ and edges $j\to i$ iff. $j\in\mathrm{pa}(i)$ is acyclic;
  2. (ii)
    a joint distribution $\mathbb{P}_{\mathbf{U}}$ over the exogenous variables.
The induced distribution $\mathbb{P}_{\mathbf{V}}$ of $\mathcal{M}$ is given by the push-forward of $\mathbb{P}_{\mathbf{U}}$ via ${\mathcal{S}}$. An intervention replaces ${\mathcal{S}}$ by a new set of assignments ${\mathcal{S}}^{\prime}$ such the graph induced by ${\mathcal{S}}^{\prime}$ is acyclic. The interventional distribution is the induced distribution $\mathcal{M}^{\prime}=({\mathcal{S}}^{\prime},\mathbb{P}_{\mathbf{U}})$. [^7] Interventions can also introduce new sources of randomness $\mathbf{U}^{\prime}$, which, for sake of simplicity, we do not consider here.
Our assumed generative process for the effect of perturbations (§ 3) is, in general, different from interventions in an SCM. However, as we show next, certain types of SCMs and interventions are recovered as a special case of our model.
Shift Interventions in Linear SCMs as a Special Case.
Consider a linear SCM with ${d_{Z}}$ endogenous variables $\mathbf{Z}$ and ${d_{Z}}$ exogenous variables $\mathbf{U}$ of the form
$\mathbf{Z}:=\mathbf{A}\mathbf{Z}+\mathbf{U},$ (C.2)
where $\mathbf{A}$ is a (lower-triangular) weighted adjacency matrix. The observational (i.e., non-intervened) distribution of $\mathbf{Z}$ induced by (C.2) is most easily understood via the reduced form expression
$\mathbf{Z}=(\mathbf{I}-\mathbf{A})^{-1}\mathbf{U}$ (C.3)
and thus given by $\mathbb{P}_{\mathbf{Z}}=(\mathbf{I}-\mathbf{A})^{-1}_{\#}\mathbb{P}_{\mathbf{U}}$. Now consider the class of shift interventions parametrised by constant shift vectors $\mathbf{c}_{e}\in\mathbb{R}^{{d_{Z}}}$, which modify the original SCM in (C.2) to
$\mathbf{Z}_{e}:=\mathbf{A}\mathbf{Z}_{e}+\mathbf{U}+\mathbf{c}_{e}.$ (C.4)
Analogous to (C.3), the reduced form of (C.4) is given by
$\mathbf{Z}_{e}$ $=(\mathbf{I}-\mathbf{A})^{-1}\mathbf{U}$ $+(\mathbf{I}-\mathbf{A})^{-1}\mathbf{c}_{e}$ (C.5)
$=\qquad\mathbf{Z}$ $+(\mathbf{I}-\mathbf{A})^{-1}\mathbf{c}_{e},$
where the second equality follows from (C.3).
Thus, if we take the shift vectors as perturbation labels (i.e., $K={d_{Z}}$ and $\mathbf{l}_{e}=\mathbf{c}_{e})$ and use a linear perturbation model of the form
$\mathbf{Z}^{\mathrm{pert}}_{e}=\bm{\phi}(\mathbf{Z}^{\mathrm{base}}_{e},\mathbf{l}_{e})=\mathbf{Z}^{\mathrm{base}}_{e}+\mathbf{W}\mathbf{l}_{e}$ (C.6)
with $\mathbf{W}=(\mathbf{I}-\mathbf{A})^{-1}$, then our perturbation model captures shift interventions in a linear SCM with adjacency matrix $\mathbf{A}$.
Remark C.2.
The above argument does not require causal sufficiency: it still holds if $\mathbb{P}_{\mathbf{U}}$ is not factorized (e.g., due to hidden confounding).
Remark C.3.
When trained on data generated according to (C.4), the adjacency matrix associated with the learnt perturbation matrix $\widehat{\mathbf{W}}$ is given by $\widehat{\mathbf{A}}=\mathbf{I}-\widehat{\mathbf{W}}^{-1}$.
Remark C.4.
The correspondence between mean shift perturbations and shift interventions appears to only hold for linear SCMs. Consider, for example, a nonlinear additive noise model,
$\mathbf{Z}:=\mathbf{f}(\mathbf{Z})+\mathbf{U},$ (C.7)
with reduced form given by
$\mathbf{Z}=\mathbf{g}(\mathbf{U}),$ (C.8)
where $\mathbf{g}$ is the inverse of the mapping $\mathbf{z}\mapsto\mathbf{z}-\mathbf{f}(\mathbf{z})$. For shift interventions in (C.7) to match our perturbation model, we then must have
$\mathbf{Z}_{e}=\mathbf{g}(\mathbf{U}+\mathbf{c}_{e})=\mathbf{g}\left(\mathbf{g}^{-1}(\mathbf{Z})+\mathbf{c}_{e}\right)=\bm{\phi}\left(\mathbf{Z},\mathbf{l}_{e}\right)$ (C.9)
for suitable $\bm{\phi}$ and $\mathbf{l}_{e}$. Thus, if $\mathbf{f}$ is nonlinear, so is $\mathbf{g}$ and therefore $\bm{\phi}$. In other words, shift interventions in a nonlinear SCM do not, in general, amount to mean shift perturbations (i.e., linear $\bm{\phi}$).
Appendix D Additional Experimental Details
Complete List of Test Cases.
The six ID and OOD test cases used for evaluation in the experiments from § 7 are listed in Tab. 2. The resulting ground truth distributions for all synthetic train and test distributions are visualized in Fig. 4.
Table 2: Test Perturbation Labels $\mathbf{l}_{\mathrm{test}}^{\mathrm{ID}}$ and $\mathbf{l}_{\mathrm{test}}^{\mathrm{OOD}}$ for the Experiment on 2D Data From § 7.
ID OOD
[1, 1, 0]11 0[1,\ 1,\ 0][ 1 , 1 , 0 ] [1, 0, 1]1 01[1,\ 0,\ 1][ 1 , 0 , 1 ]
[0.1, 0.25, 0]0.10.25 0[0.1,\ 0.25,\ 0][ 0.1 , 0.25 , 0 ] [0, 1, 1]011[0,\ 1,\ 1][ 0 , 1 , 1 ]
[0.5, 0.75, 0]0.50.75 0[0.5,\ 0.75,\ 0][ 0.5 , 0.75 , 0 ] [0, 2, 0]02 0[0,\ 2,\ 0][ 0 , 2 , 0 ]
[0.5, 0.5, 0]0.50.5 0[0.5,\ 0.5,\ 0][ 0.5 , 0.5 , 0 ] [2, 0, 0]2 0 0[2,\ 0,\ 0][ 2 , 0 , 0 ]
[0, 0.5, 0]00.5 0[0,\ 0.5,\ 0][ 0 , 0.5 , 0 ] [0.5, 1.5, 0]0.51.5 0[0.5,\ 1.5,\ 0][ 0.5 , 1.5 , 0 ]
[0.5, 0, 0]0.5 0 0[0.5,\ 0,\ 0][ 0.5 , 0 , 0 ] [1.5, 0.5, 0]1.50.5 0[1.5,\ 0.5,\ 0][ 1.5 , 0.5 , 0 ]
Figure 4: Complete Ground Truth Training and Test Data. Shown are all train and test distributions, with crosses marking the respective means. The training cases are depicted in grey, the ID test cases in blue and the OOD test cases in orange.
Hyperparameter Selection for CPA.
To improve the fit of CPA on the synthetic dataset, we perform a grid search over key hyperparameters. Specifically, we consider learning rates of the autoencoder and adversarial classifier from {1e-2, 5e-3, 1e-3, 5e-4, 1e-4, 5e-5}, and a latent bottleneck dimension from $\hat{d_{Z}}$ in {2, 4, 8}. We select the best joint parameter configuration in terms of a combined performance measure (ED + MMD + mean difference) on a held out portion of 25% of the original training data, which yields a learning rate of 1e-3 for the autoencoder, a learning rate of 5e-3 for the adversarial classifier, and a bottleneck dimension of 4. All other hyperparameters are set to their default values.
Appendix E Additional Results
In this experiment, each observation $\mathbf{x}_{e,i}$ is generated from two signal dimensions $\mathbf{z}_{e,i}^{\text{pert}}\in\mathbb{R}^{2}$ and eight noise dimensions $\varepsilon_{e,i}^{\text{pert}}\in\mathbb{R}^{8}$ as input for the nonlinear mixing function, as described at the end of § 7. We evaluate model robustness under increasing levels of noise by varying the standard deviation $\sigma_{\varepsilon}$ of the noise variable $\varepsilon\sim\mathbb{Q}_{\epsilon}$ in the data-generating process. The list of values of $\sigma_{\epsilon}$ considered is [0, 0.01, 0.1, 0.25, 0.5, 1, 2, 10]00.010.10.250.51210[0,\ 0.01,\ 0.1,\ 0.25,\ 0.5,\ 1,\ 2,\ 10][ 0 , 0.01 , 0.1 , 0.25 , 0.5 , 1 , 2 , 10 ]. Both PDAE and CPA models are trained for 2000 epochs. One random seed is used and shared across data generation, model initialization, training, and evaluation. We use the same hyperparameters as described in § 7, with the exception of PDAE, for which we increase the reconstruction loss weight $\lambda=0.25$ and set the decoder noise dimension to 1 with a standard deviation of 0.5; and CPA, for which we use a 2-dimensional embedding. The results are summarised in Fig. 5. As can be seen, PDAE performs competitively across a range of $\sigma_{\epsilon}$ (especially low to moderate ones), despite only having access to one noise dimension in the decoder. As $\sigma_{\epsilon}$ grows large, most signal is lost, and the performance of all methods tends to that of the pooling baseline, as expected.
Figure 5: Evaluation of Robustness Under Increasing Levels of Noise. We compare the ID test performance of the baseline methods, CPA, and PDAE for higher dimensional observations with additional noise dimensions appended while increasing the level of noise $\sigma_{\varepsilon}$.